OpenStudy (anonymous):

Solve simultaneously equations 13) x/3 - y/4 = 5 X/2. + y/3 = -1

OpenStudy (lulubj):

this is history?

OpenStudy (anonymous):

Go to mathematics and post that.

OpenStudy (texaschic101):

x/3 - y/4 = 5 (multiply by the LCD which is 12) 4x - 3y = 60 x/2 + y/3 = -1 (multiply by the LCD which is 6) 3x + 2y = -6 now we have : 4x - 3y = 60 -->(2) 3x + 2y = -6 -->(3) ---------------- 8x - 6y = 120 (result of multiplying by 2) 9x + 6y = -18 (result of multiplying by 3) ----------------add 17x = 102 x = 102/17 x = 6 now sub 6 in for x in either equation 3x + 2y = -6 3(6) + 2y = -6 18 + 2y = -6 2y = -6 - 18 2y = - 24 y = -12 check... x/2 + y/3 = -1 6/2 + (-12/3) = -1 3 - 4 = -1 -1 = -1 (correct) x/3 - y/4 = 5 6/3 -(-12/4) = 5 2 + 3 = 5 5 = 5 (correct) x = 6 and y = -12

OpenStudy (anonymous):

You should be glad you have someone smart in Mathematics here!

OpenStudy (texaschic101):

Thank you Kfins :)